OpenStudy (anonymous):

How do I figure out what sin(theta - pi) equals? The answer choices are (a) cos(theta) (b) -sin(theta) or (c) sin(theta)

myininaya (myininaya):

Do you know the difference formula for sine?

OpenStudy (helder_edwin):

try this \[ \large \sin(A+B)=\sin A\cdot\cos B+\sin B\cdot\cos A \] and remember that \[ \large \sin\pi=0 \]

myininaya (myininaya):

sum formula works to since any difference can be written as a sum

OpenStudy (helder_edwin):

right. \[ \large \sin(A-B)=\sin A\cdot\cos B-\sin B\cdot\cos A \] sorry about the typo.